OpenStudy (anonymous):

Solve. x2 + 4x + 1 = 0

OpenStudy (theviper):

Use quadratic formula:-\[\Huge{\color{blue}{-b \sqrt{b^2-4ac}\over2a}}\]

OpenStudy (theviper):

a=1, b=4, c=1 Put the values & get ur answer:)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

-4\[-4\sqrt{b ^{2}}-4(1)(1)/2(1)\]

OpenStudy (campbell_st):

you need to general quadratic formula or you need to complete the square the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 1, b = 4 and c = 1 or by completing the square let \[x^2 + 4x + 1 = 0\] find \[(\frac{4}{2})^2 = 4\] you need the constant term to be 4 so add 3 to both sides then \[x^2 + 4x + 4 = 3\] factorise \[(x + 2)^2 = 3\] take the square root of both sides \[x + 2 = \pm \sqrt{3}\] subtract 2 from both sides \[x = -2 \pm \sqrt{3}\]